A normal distribution has μ = 55 and σ = 5. What is the probability of randomly selecting a score greater than or equal to 55 from this distribution?

a. About 16%

b. About 84%

c. About 20%

d. About 22%
Answer #1

z score for 55 = (55 - 50)/5 = 1

Hence,

P(Selecting a score greater than or equal to 55 from this distribution)

= P(z > 1)

= About 16%

Option A is correct.